Solve each system of equations:
2x= y+12
3x+2y= -3
What I did:
(2x=y+12)2
3x=-2y-3
4x=2y+28
3x=-2y-3

Found 2 solutions by TimothyLamb, Alan3354:
Answer by TimothyLamb(4379)

---
2x = y + 12
3x + 2y = -3
---
put the system of linear equations into standard form:
---
2x - y = 12
3x + 2y = -3
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 3
y = -6
---
system matrix
2 -1
3 2
---
inverse of system matrix
0.28571429 0.14285714
-0.42857143 0.28571429
---
determinant of system matrix = 7
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by Alan3354(69443)

Solve each system of equations:
2x= y+12
3x+2y= -3
What I did:
(2x=y+12)2
3x=-2y-3
4x=2y+28
3x=-2y-3
=============
What I would do:
2x= y+12 --> 2x - y = 12
3x+2y= -3
-------
Put in standard form
2x- y = 12
3x +2y= -3
-----
Multiply eqn 1 by 2
4x- 2y = 24
3x +2y= -3
-------------------- Add
7x = 21
x = 3
-----
Sub for x in any eqn and find y